In the manufacturing of a certain article, two types of defects are noted that occur with probabilities 0.02, and 0.05, respectively. If these two defects occur independently of one another what is the probability that an article is free of both kinds of defects?
Answer #1

Solution

P(A) = 0.02, P(B) =0.05

P(A') = 0.98, P(B') = 0.95

So P(A' intersection B') = P(A') * P(B') = 0.98*0.95 = 0.931

So there is 93.1 % chances that an article is free or both kind of defects.